OpenStudy (anonymous):

medal and fan The numbers 20, 21, and c are a Pythagorean triple. What is the value of c? A. 22 B. 29 C. 41 D. 25

OpenStudy (anonymous):

I really don't know how to explain this, what do you think it is?

OpenStudy (anonymous):

idk thats why i posted the question ..

OpenStudy (anonymous):

Right... http://nrich.maths.org/1309 says, "These...give the Pythagorean triple a=20, b = 21 and c=29" So the last part of the triple is 29.

OpenStudy (anonymous):

thankyou :)

jimthompson5910 (jim_thompson5910):

solve for c a^2 + b^2 = c^2 20^2 + 21^2 = c^2

OpenStudy (anonymous):

You could also use that ^